OpenStudy (anonymous):

Find the sum a1 = 625, n= 5, r = 3/5 a1 = 80, n= 7, r = - 1/2 a1 = 343, an = -1, r = -1/7 @campbell_st @robtobey

OpenStudy (campbell_st):

ok... so these are all geometric series... where the common ratio is between -1 < r < 1 this means that each series had a limiting sum.... which is found using \[S_{\infty} = \frac{a_{1}}{1 - r}\] just substitute and evaluate the alternate method , given you have a value for n, the number of terms is \[S_{n} = \frac{a_{1}(1 - r^n)}{1 - r}\] just substitute and calculate hope it helps